OpenStudy (anonymous):

solve. sqrt 20 + sqrt 40

OpenStudy (anonymous):

\[\sqrt{4*5}+\sqrt{4*10}=2\sqrt{5}+2\sqrt{10}\]

OpenStudy (anonymous):

thank you!